FEL Upgrade Project Weekly Brief –May 20-24, 2002




The wiggler for the 10 kW Upgrade was installed on its mounting bolts in the final position in the FEL vault this week. There was also considerable progress made in the installation of stands for magnets in the recirculation arcs and with the support stands for the second and third linac cryomodules.

A critical final test run for ion implantation of the Upgrade gun cathode support tube was successfully completed this week. At week's end the first of six runs to implant the actual support tube was also successfully completed.


The FEL project management team (Dylla and Neil) attended a meeting of the FEL technical Working Group (FEL-TAWG) on Wednesday May 22 sponsored by the DoD Joint Technology Office. The TAWG is being re-constituted under a new chair Michael Dietchman, from ONR , who is taking over after the retirement of Eli Zimet in January. Many of the FEL experts that have been serving on the FEL Advisory Committee for the FEL Upgrade project have been invited to be members of the TAWG. The purpose of this working group is to advise the JTO on FEL R&D.

On Thursday, May 23 we attended a workshop in Washington on directed energy activities sponsored by the Directed Energy Professional Society (DEPS). The Navy's interest and ownership of the FEL development was presented by RADM Mathis and Cmd. Roger McGinnis of NAVSEA. The JLab FEL program was given good press at this meeting by both the above speakers and by the secretary of last year's Defense Science Board panel on high energy laser development.

DOE approved the JLab CRADA with Advanced Energy Systems for the development of a high current injector (100 mA) to support the 100 kW Upgrade program.

F. Dylla presented a talk on applications of the JLab FEL to surface analysis and surface processing to the AVS sponsored Topical Conference on Surface Analysis held on May 21-23 at Vanderbilt University.




WBS 4 (Injector):

Continued to work with W&M on the ion implantation system for the gun electrodes. Three implantation runs performed on a fourth test tube were completed successfully, meaning that the problems of arcing and implanted coating peeling off have been resolved. The actual support tube was mounted yesterday and the first of six runs has been completed successfully. Continued polishing the ball cathode which will be implanted next. Worked on the designs for the SF6 tank modifications and the HV resistor.

Gun HVPS - The 1st conditioning resistor for the Gun HVPS was sprayed and is drying. The 2nd and 3rd will be sprayed the next 2 weeks after which all 3 will be fired and tested. The tank for the HVPS is now due 6/6/02.

WBS 6 (RF):

Quarter HVPS - Completed low level wiring of unit 3 and started testing Heater Voltage circuits. Heater Voltage board for unit 4 is finished, but not yet installed. The PLC code is not expected until 3rd week of June.

Quarter Klystrons - The parts are finished for the 2nd klystron cart, but they are not yet completely assembled. The parts for the 3rd klystron cart are in progress. Both are expected to be assembled within 2 weeks.



WBS 8 (Instrumentation):

Work continued on fabricating additional OMS chassis'. The lens mirror covers for the beamviewers were painted and ready for installation onto the frames. Progress was made on the cable pull for the Ion Pumps. The new smoke detectors were interlocked into the LSS. The PLCs for the MPS, CAMAC crate and modules for the BLMs have been removed from the racks along with a clean up of the cables.

FL06B01 (in the drive laser clean room) has been upgraded to include a VME crate running a new ioc (iocfel10) for the new drive laser pulse control system. This effort included a reorganization of existing hardware and a re-working of the cabling in the rack. Work continues on the Inventory Control System and significant improvements have been posted to the server. The first attempt at a check print for the BLM vme card (F0151) is being reviewed and appended.

WBS 9 (Transport):


Optical Chicane Dipoles (GW)

• At Magnet Enterprises International (MEI) in Oakland CA. We are studying the epoxy problems.

• Upper and lower core’s plates remain in final machining. We will check completed cores on June 7

Injector Dipoles (DU/DV)

• At Master Machine, core pieces remain in final machining.

• At Wang NMR, they remain working on potting the coils.

Arc Dipoles (GY, GX, GQ)

• Process Equipment Co. (PECo) continued manufacturing the cores and parts of the GQ-GX dipoles. GQ cores are at heat treat. GX cores are in rough machining

• Wang NMR continues working on the winding fixture for the GQ-GX coils.

• AES drew up the shim system for the use of these magnets at the highest energy. The drawings were checked and sent to PECO for incorporation into their production.

• Wang NMR started winding the second GY coil and is working on the potting fixture.

• Bosma Machine is shipping the first article GY Core to WANG for assembly. We will inspect the second core’s gap on June 5th.


3 inch quad (QX)

• We found that the electric current set point reliability of the Trim Card Power Supplies is about ±3 parts in 1000, well below David Douglas’ specification for magnetic field set ability of 1 part per 1000. Grist for getting better trim supplies. After some down time on the magnet test stand, the full production retesting of the QX quads started.

Trim Quad (QT)

• Milhous Control of Virginia has almost all coils made so he can deliver all the remaining magnets in one lot.

• No word on the GC corrector because Tom Hiatt is consumed with test stand problems.

Sextupole (SF)

• DULY Research continues to work on completing drawings of the magnet. They are finishing up details.

Octupole (OT)

• Work on this magnet is on furlough until the sextupole is designed.

Beam Line and Vacuum

• The procurement documents for the ARC Chambers are in composition.

• A number of stands were delivered and bolted into position.

• Work continues on the recirculation region in front of the Cryomodule.

• Design of the optical chicane region with its chamber continues.

• Design of the regions around the wiggler continues.

WBS 10 (Wiggler):

The wiggler was installed onto its stands this week and is now ready for alignment and diagnostic installation.

WBS 11 (Optics):

Detailing the drawings for the optical cavity vacuum vessels has been somewhat more time-consuming than originally projected, so the plan is to release preliminary drawings to the fabrication scheduling team next week, and final drawings the week after. We met and discussed the request for fabrication advice by the vendor of the optical cavity internals. We determined that the SOW contains sufficient guidance, and signed off on the vendor's plan to machine both sides of all plates.

We held a planning meeting to work out the assembly and test plan for these parts. A separate meeting was held with the fabrication scheduler to have him determine that all fasteners, etc, will be available to us when parts begin appearing in several weeks.

At the time of this writing we are awaiting word on the delivery of one OCMMS assembly. The other OCMMS assembly continues to be delayed because of late delivery of the calcium fluoride vacuum viewports. We received the external OCMMS mirrors from the vendor 3 wks early. They will be inspected and measured to see that they meet specs.

Fabrication the first article 50kW optical beam dump is complete, except for brazing. This will be done within the next two weeks. The second "foam metal" (a metal mesh) design is also complete, but requires the metal mesh before it can be brazed. Vendors to create the absorber on the copper were identified and will be contacted to get pricing and delivery.

Our additional laser beam profiler arrived and was installed. We met with the software engineers to iterate plans for getting the new IR beam profiler output into EPICS.

We were informed by the vendor of the ultrafast laser oscillator that it will not ship before today. Safety goggles and glasses for this laser system were received. Assuming the safety audit proceeds on schedule, we are on track to have it installed the week of June 3.

The THz radiation monitor mechanical design was reviewed with respect to minimizing wakefields. Solutions were proposed (adding a trapdoor) and this is being incorporated into the design.

Other activities:

We are looking at ways to induce chirp on the electron beam (and thus the FEL output) in such a way that the electron beam can still be energy recovered. This will be needed when we wish to compress the pulsewidth of the FEL.